As a settlement for an insurance claim, Craig was offered one of two choices. He could either accept a lump-sum amount of

$10,307

now, or accept

monthly

payments of

$144

for the next

seven

years. If the money is placed into a trust fund earning

4.92 %

compounded

semi-annually ,

which is the better option and by how much?
Answer #1

Option 1: Lumpsum Amount = $ 10307

Option 2: Monthly Payments worth $ 144, Tenure = 7 years or (7 x 12) = 84 months, Annual Interest Rate = 4.92 %, Compounding Frequency: Semi-Annual

Semi-Annual Rate = 4.92 / 2 = 2.46 %

Equivalent Monthly Rate = (1.0246)^(1/6) - 1 = 0.0040586 or 0.40586 %

Therefore, Present Value of Monthly Payments = 144 x (1/0.0040586) x [1-{1/(1.0040586)^(84)}] = $ 10232.3216 ~ $ 10232.32

As the lumpsum amount pays more, Option 1 should be selected.

Additional Amount Received = 10307 - 10232.32 = $ 74.6784 ~ $ 74.68